# Vienna Mosaic: Navigating Social Borders in a Melting Pot

### v1

@@ Methods
To support our analyses of the residential patterns of inhabitants in Vienna, we used several additional datasets. The geographic boundaries for the Vienna districts were obtained from the Vienna GIS database [45], while the global GIS data provided country boundaries to facilitate spatial visualisation and cross-national comparisons [46]. Population counts by district were sourced from census data filtering for the year 2023, allowing normalisation of migrant fraction within districts and the extraction of the Austrian national population [47]. The average net income data at the district level were filtered for the year 2020 and served as a proxy for economic status, helping to assess the socioeconomic factors that influence residential clustering [48]. Additional datasets are detailed in the Supplementary Note LABEL:sec:otherdatasets.
@@
In this section, we outline the formalisation of our model for measuring the co-residence network between nationalities in the Vienna districts. This network model allows us to identify significant clustering patterns within Vienna’s spatial structure. The connections, or links, between nationalities are based on the degree to which two nationalities co-reside in the same districts [32]. To compute co-residence links between countries, self-loops, representing co-residence within the same nationality, are excluded from the analysis. This ensures that the resulting network focuses solely on interactions and shared residential patterns between different nationalities, allowing for a clearer investigation of cross-national co-residence dynamics.
@@
Given that observed link weights may vary widely due to factors such as population size and district-level density, it is necessary to determine the statistical significance of these links. To do so, we compare the empirically observed link weights to a null model. In the null model, we assume that each resident comes from a country randomly picked from the total population of Vienna. The random assignments are performed by drawing the countries’ population within each district from a multinomial distribution. This random distribution accounts for the baseline expectation that two nationalities might co-reside by chance alone, based on their overall presence in the city. From the null model, the analytical expected probability $\mu_{ij}^{d}$ that two nationalities $i$ and $j$ co-reside in a given district $d$ can be expressed as [32]:
@@
where $\sigma_{ij}^{d}$ is the standard deviation of the null distribution for link weights. Using the multinomial theorem multiple times, one can compute the standard deviation as [32]:
@@
To find these clusters, we employ a community detection approach that uses the strength of co-residence connections between nationalities across Vienna’s districts as a basis for grouping. The network of co-residence is constructed with each nationality pair $(i,j)$ linked by their co-residence weights $z_{ij}$, which reflect the frequency of their spatial overlap across districts (as detailed in the previous subsection and shown in Fig. 2b). This network provides the structural foundation to reveal groups of countries with significant shared residence patterns. Our approach to clustering relies on the network community-detection method known as the map equation, which leverages random walks to detect communities based on their flow within the network [37, 38]. In this context, the map equation algorithm simulates a sequence of “visits” to nationalities, where movement between nationalities is determined by the strength of their co-residence link. This method clusters nationalities based on how often a random walker would transition between them, prioritising strongly interconnected groups that are frequently co-resided within the same districts. Nationalities with stronger internal connections are likely to retain the walker for extended periods, which results in distinct, identifiable clusters. Here, we use the map equation’s associated search algorithm Infomap [39] to identify the groups of countries.
@@
Here, we leverage our dataset of migrants and locals in Vienna to investigate the inherent diversity within the city’s districts, as well as the degree of country-based homophily among residents. To achieve this, we apply two distinct statistical indices: the Simpson index and the Dissimilarity index. The Simpson index is used to quantify the diversity within each district, providing insight into the overall heterogeneity of the resident population in terms of nationality. A higher Simpson index reflects a more diverse population within a district, indicating a greater variety of nationalities living in close proximity. On the other hand, the Dissimilarity index is employed to measure homophily, or the tendency of individuals from the same nationality to cluster together within certain districts. This index helps us assess how much the population of each nationality tends to concentrate in specific areas rather than being evenly distributed across the city [49, 50]. By analysing both the Simpson index for district diversity (see Fig. 4a and Fig. 4b) and the Dissimilarity index for homophily (see Fig. 4c and Fig. 4d), we gain valuable insights into the underlying reasons why certain nationalities in Vienna segregate into the two clusters identified in our analysis.
The Simpson index, which can be calculated for each district (local) and the entire city, is a measure of concentration. It expresses the probability that two randomly selected individuals from the district/city have different nationalities. Calculated for one district d of the city, the Simpson index, $S^{d}$, becomes [51, 52]:
@@
where $P_{i}^{d}=N_{i}^{d}/N^{d}$ is the fraction of inhabitants of nationality i residing in district d. It is a measure between 0 and a maximum value, which depends on the different number of nationalities, $C^{d}$, living in the district. The higher $S^{d}$, the more nationality-diverse the district is. A district is fully diverse ($S^{d}_{\text{max}}=1-1/C^{d}$) when the population in the district $d$ is equally distributed among all existing nationalities. In contrast, a district has no diversity ($S^{d}_{\text{min}}=0$), when it is inhabited exclusively by a single nationality. The Simpson index exhibits a close mathematical relationship with entropy, which serves as a quantitative measure of disorder (see Supplementary Note LABEL:sec:entropy). This relationship is often utilised within segregation studies to assess the unequal distribution of national groups across neighbourhoods or between global and local scales [53].
Furthermore, the city-wide Simpson index can be derived analogously to Eq. (8) by using the full population distribution in Vienna. The value for the city of Vienna is $S_{\text{City}}$ = 0.541. This number serves as a reference to compare with the average local Simpson index, which is the population-weighted average of the local Simpson index values in all districts. It is computed as [43]:
@@
To define national homophily, the tendency to reside among others of the same nationality i, we calculate the Dissimilarity index $D_{i}$, which indicates the percentage of members of that group who would need to relocate so that the distributions at the district-level match the distribution at the city-level [43]:
@@
The Pearson correlation coefficient, denoted as r, measures the strength and direction of a linear relationship between two continuous variables. It ranges from -1 to +1, where $r=+1$ indicates a perfect positive linear relationship, $r=-1$ denotes a perfect negative linear relationship, and $r=0$ suggests no linear association between the variables [54]. Interpretation of r typically follows Cohen’s guidelines, which suggest that $|r|<0.3$ represents a weak relationship, $0.50.3\leq|r|<$ 0.5 indicates a moderate relationship, and $|r|\geq 0.5$ suggests a strong relationship [55]. However, these cutoffs are context-dependent, and scientific judgment should consider the practical significance of the relationship in the specific field of study.
@@
where $\bar{X}$ and $\bar{Y}$ are the sample means of X and Y, respectively. This equation standardises the covariation of X and Y by the product of their standard deviations, yielding a dimensionless value that facilitates comparison across different datasets and units [56].
@@
Pearson correlations shown in Fig. 3 were calculated for both subpanels a and b using the population fraction of each nationality relative to the total district population (see Supplementary Fig. LABEL:fig:figs4). In subpanel a, this population fraction was correlated with the average net income per district, derived from the additional dataset on income in Vienna’s districts [48] (see Supplementary Fig. LABEL:fig:figs2a). In subpanel b, the correlations were computed using the difference between each district’s diversity and the city-wide diversity (see Fig. 4b). Detailed Pearson correlation results for income and diversity are provided in Supplementary Table LABEL:tab:tabs4.
@@
In GitHub repository [36], we provide the aggregate statistics on the proportion of citizens of each nationality living in each district. Together with the publicly available complementary datasets, these data are sufficient to reproduce all the analyses performed in this study.
@@
The code developed to perform all analyses is available in GitHub repository [36].

### v2

@@ Methods
To support our analyses of the residential patterns of inhabitants in Vienna, we used several additional datasets. The geographic boundaries for the Vienna districts were obtained from the Vienna GIS database (City of Vienna, 2023c, ), while the global GIS data provided country boundaries to facilitate spatial visualisation and cross-national comparisons (Natural Earth,, 2023). Population counts by district were sourced from census data filtering for the year 2023, allowing normalisation of migrant fraction within districts and the extraction of the Austrian national population (City of Vienna, 2023b, ). The average net income data at the district level were filtered for the year 2020 and served as a proxy for economic status, helping to assess the socio-economic factors that influence residential clustering (City of Vienna, 2023a, ). Additional datasets are detailed in the Supplementary Note LABEL:sec:otherdatasets.
@@
In this section, we outline the formalisation of our model for measuring the co-residence network between nationalities in the Vienna districts. This network model allows us to identify significant clustering patterns within Vienna’s spatial structure. The connections, or links, between nationalities are based on the degree to which two nationalities co-reside in the same districts (Karimi et al.,, 2015). To compute co-residence links between countries, self-loops, representing co-residence within the same nationality, are excluded from the analysis. This ensures that the resulting network focuses solely on interactions and shared residential patterns between different nationalities, allowing for a clearer investigation of cross-national co-residence dynamics.
@@
Given that observed link weights may vary widely due to factors such as population size and district-level density, it is necessary to determine the statistical significance of these links. To do so, we compare the empirically observed link weights to a null model. In the null model, we assume that each resident comes from a country randomly picked from the total population of Vienna. The random assignments are performed by drawing the countries’ population within each district from a multinomial distribution. This random distribution accounts for the baseline expectation that two nationalities might co-reside by chance alone, based on their overall presence in the city. From the null model, the analytical expected probability $\mu_{ij}^{d}$ that two nationalities $i$ and $j$ co-reside in a given district $d$ can be expressed as (Karimi et al.,, 2015):
@@
where $\sigma_{ij}^{d}$ is the standard deviation of the null distribution for link weights. Using the multinomial theorem multiple times, one can compute the standard deviation as (Karimi et al.,, 2015):
@@
To find these clusters, we employ a community detection approach that uses the strength of co-residence connections between nationalities across Vienna’s districts as a basis for grouping. The network of co-residence is constructed with each nationality pair $(i,j)$ linked by their co-residence weights $z_{ij}$, which reflect the frequency of their spatial overlap across districts (as detailed in the previous subsection and shown in Fig. 2b). This network provides the structural foundation to reveal groups of countries with significant shared residence patterns. Our approach to clustering relies on the network community-detection method known as the map equation, which leverages random walks to detect communities based on their flow within the network (Rosvall and Bergstrom,, 2008; Rosvall et al.,, 2009). In this context, the map equation algorithm simulates a sequence of “visits” to nationalities, where movement between nationalities is determined by the strength of their co-residence link. This method clusters nationalities based on how often a random walker would transition between them, prioritising strongly interconnected groups that are frequently co-resided within the same districts. Nationalities with stronger internal connections are likely to retain the walker for extended periods, which results in distinct, identifiable clusters. Here, we use the map equation’s associated search algorithm Infomap (Edler et al.,, 2023) to identify the groups of countries.
@@
Here, we leverage our dataset of migrants and locals in Vienna to investigate the inherent diversity within the city’s districts, as well as the degree of country-based homophily among residents. To achieve this, we apply two distinct statistical indices: the Simpson index and the Dissimilarity index. The Simpson index is used to quantify the diversity within each district, providing insight into the overall heterogeneity of the resident population in terms of nationality. A higher Simpson index reflects a more diverse population within a district, indicating a greater variety of nationalities living in close proximity. On the other hand, the Dissimilarity index is employed to measure homophily, or the tendency of individuals from the same nationality to cluster together within certain districts. This index helps us assess how much the population of each nationality tends to concentrate in specific areas rather than being evenly distributed across the city (Simpson,, 2007; Harris,, 2017). By analysing both the Simpson index for district diversity (see Fig. 4a and Fig. 4b) and the Dissimilarity index for homophily (see Fig. 4c and Fig. 4d), we gain valuable insights into the underlying reasons why certain nationalities in Vienna segregate into the two clusters identified in our analysis.
The Simpson index, which can be calculated for each district (local) and the entire city, is a measure of concentration. It expresses the probability that two randomly selected individuals from the district/city have different nationalities. Calculated for one district d of the city, the Simpson index, $S^{d}$, becomes (Duncan and Duncan,, 1955; White,, 1986):
@@
where $P_{i}^{d}=N_{i}^{d}/N^{d}$ is the fraction of inhabitants of nationality i residing in district d. It is a measure between 0 and a maximum value, which depends on the different number of nationalities, $C^{d}$, living in the district. The higher $S^{d}$, the more nationality-diverse the district is. A district is fully diverse ($S^{d}_{\text{max}}=1-1/C^{d}$) when the population in the district $d$ is equally distributed among all existing nationalities. In contrast, a district has no diversity ($S^{d}_{\text{min}}=0$), when it is inhabited exclusively by a single nationality. The Simpson index exhibits a close mathematical relationship with entropy, which serves as a quantitative measure of disorder (see Supplementary Note LABEL:sec:entropy). This relationship is often utilised within segregation studies to assess the unequal distribution of national groups across neighbourhoods or between global and local scales (Finney and Simpson,, 2009).
Furthermore, the city-wide Simpson index can be derived analogously to Eq. (8) by using the full population distribution in Vienna. The value for the city of Vienna is $S_{\text{City}}$ = 0.541. This number serves as a reference to compare with the average local Simpson index, which is the population-weighted average of the local Simpson index values in all districts. It is computed as (Zuccotti et al.,, 2023):
@@
To define national homophily, the tendency to reside among others of the same nationality i, we calculate the Dissimilarity index $D_{i}$, which indicates the percentage of members of that group who would need to relocate so that the distributions at the district-level match the distribution at the city-level (Zuccotti et al.,, 2023):
@@
The Pearson correlation coefficient, denoted as r, measures the strength and direction of a linear relationship between two continuous variables. It ranges from -1 to +1, where $r=+1$ indicates a perfect positive linear relationship, $r=-1$ denotes a perfect negative linear relationship, and $r=0$ suggests no linear association between the variables (Pearson,, 1895). Interpretation of r typically follows Cohen’s guidelines, which suggest that $|r|<0.3$ represents a weak relationship, $0.50.3\leq|r|<$ 0.5 indicates a moderate relationship, and $|r|\geq 0.5$ suggests a strong relationship (Cohen,, 2013). However, these cutoffs are context-dependent, and scientific judgment should consider the practical significance of the relationship in the specific field of study.
@@
where $\bar{X}$ and $\bar{Y}$ are the sample means of X and Y, respectively. This equation standardises the covariation of X and Y by the product of their standard deviations, yielding a dimensionless value that facilitates comparison across different datasets and units (Fisher,, 1970).
@@
Pearson correlations shown in Fig. 3 were calculated for both subpanels a and b using the population fraction of each nationality relative to the total district population (see Supplementary Fig. LABEL:fig:figs4). In subpanel a, this population fraction was correlated with the average net income per district, derived from the additional dataset on income in Vienna’s districts (City of Vienna, 2023a, ) (see Supplementary Fig. LABEL:fig:figs2a). In subpanel b, the correlations were computed using the difference between each district’s diversity and the city-wide diversity (see Fig. 4b). Detailed Pearson correlation results for income and diversity are provided in Supplementary Table LABEL:tab:tabs4.
@@
In GitHub repository (Parera,, 2024), we provide the aggregate statistics on the proportion of citizens of each nationality living in each district. Together with the publicly available complementary datasets, these data are sufficient to reproduce all the analyses performed in this study.
@@
The code developed to perform all analyses is available in GitHub repository (Parera,, 2024).
@@ Acknowledgments
Ethical approval
This article does not contain any studies with human participants performed by any of the authors.
Informed consent
This article does not contain any studies with human participants performed by any of the authors.